# Temporal Causal Discovery in Dynamic Bayesian Networks Using Federated Learning

### v1

@@ 6 Discussion
We proposed a federated framework for learning DBNs on horizontally partitioned data. Specifically, we designed a distributed DBN learning method using ADMM, where only model parameters are exchanged during optimization. Our experimental results demonstrate that this approach outperforms other state-of-the-art methods, particularly in scenarios with numerous clients, each possessing a small sample size — a common situation in federated learning that motivates client collaboration. Below, we address some limitations of our approach and suggest potential directions for future research.
@@
We assumed that the structure of the DBN is fixed over time and is identical for all time series in the dataset (i.e., it is the same for all $m\in M$). Relaxing these assumptions could be useful in various ways, such as allowing the structure to change smoothly over time [44]. Another direction for future work is to investigate the behavior of the algorithm on non-stationary or cointegrated time series [28] or in scenarios with confounding variables [17].
@@
The proposed ADMM-based approach to federated learning relies on a stateful setup, requiring each participating client to be involved in every round of communication and optimization. This “always-on” requirement can be burdensome in real-world scenarios. For instance, in large-scale deployments, clients such as mobile devices or IoT sensors may experience intermittent connectivity, limited power, or varying levels of availability. Ensuring that all such devices participate consistently and synchronously in every round is often impractical and can result in significant performance bottlenecks. An important future direction is to explore asynchronous techniques, which would enable stateless clients and facilitate cross-device learning. Furthermore, data may be vertically partitioned across clients, meaning that each client owns different variables but collectively aims to perform Bayesian Network Structure Learning (BNSL). Developing a federated approach tailored to this vertical setting represents another promising research direction. Additionally, the ADMM procedure involves sharing model parameters with a central server, raising concerns about potential privacy risks. Research has shown that these parameters can leak sensitive information in certain scenarios, such as with image data [40]. To address this, exploring differential privacy techniques [10] to enhance the protection of shared model parameters is a critical avenue for future work. Moreover, our current work focuses on non-heterogeneous data. However, heterogeneous settings are more applicable to many real-world problems, as clients often differ in computational capabilities, communication bandwidth, and local data distributions. These variations pose significant challenges for model convergence and performance. Several methods have been proposed to address these issues. For example, FedProx [24] incorporates a proximal term to stabilize and unify updates from clients with varying local training conditions, while FedNova [50] normalizes aggregated updates to mitigate the impact of heterogeneous local computational workloads.

### v2

@@ 6 Discussion
We proposed a federated framework for learning Dynamic Bayesian Networks (DBNs) on homogeneous time series data and a personalized federated framework for heterogeneous time series data. Specifically, we designed a federated DBN learning method using ADMM, where only model parameters are exchanged during optimization. Our experimental results demonstrate that this approach outperforms other state-of-the-art methods, particularly in scenarios with numerous clients, each possessing a small sample size—a common situation in federated learning that motivates client collaboration. Additionally, we developed a personalized federated DBN learning method by incorporating a proximal operator and utilizing ADMM. Our experimental results indicate that the personalized approach improves accuracy while maintaining robust performance even with small sample sizes and a large number of clients. Furthermore, we demonstrate that partial client participation does not lead to a loss of accuracy. Below, we address some limitations of our approach and suggest potential directions for future research.
@@
We assume that the DBN structure, represented by $W,A$, remains constant over time and is the same for all time series in the dataset. Relaxing this assumption could be useful in various ways, such as allowing the structure to change smoothly over time [52]. Another direction for future work is to investigate the behavior of the algorithm on non-stationary or cointegrated time series [32] or in scenarios with confounding variables [21].
@@
In our FDBNL, the ADMM-based approach to federated learning relies on a stateful setup, requiring each participating client to be involved in every round of communication and optimization. Even though we attempt to implement partial client participation for PFDBNL, applying the same strategy is meaningless since it only affects which clients join, without altering the fundamental process. This “always-on” requirement can be burdensome in real-world scenarios. For instance, in large-scale deployments, clients such as mobile devices or IoT sensors may experience intermittent connectivity, limited power, or varying levels of availability. Ensuring that all such devices participate consistently and synchronously in every round is often impractical and can result in significant performance bottlenecks. An important future direction is to apply other federated optimization techniques and explore asynchronous techniques, which would enable stateless clients and facilitate cross-device learning. Additionally, the ADMM procedure involves sharing model parameters with a central server, raising concerns about potential privacy risks. Research has shown that these parameters can leak sensitive information in certain scenarios, such as with image data [46]. To address this, exploring differential privacy techniques [12] to enhance the protection of shared model parameters is a critical avenue for future work.
Personalized Learning.
Heterogeneous settings are more applicable to many real-world problems, as clients often differ in computational capabilities, communication bandwidth, and local data distributions. These variations pose significant challenges for model convergence and performance. Thus, we propose the first methodology for learning Personalized DBNs by using the proximal operator. This framework can be easily extended to non-temporal cases, such as NOTEARS [64]. As mentioned earlier, many potential personalized federated learning frameworks can be applied, such as meta-learning and clustered federated learning. Moreover, there are several interesting research directions to explore. In our paper, we study the personalized setting under a fixed $d$, i.e., the same number of variables for each client. A promising future direction is to address scenarios where $d$ varies across different clients. Similarly, for $p$, each client may have different time-lagged dependencies they wish to explore. Additionally, the case of mixed connectivity levels is an important research direction. For instance, in the presence of sparse networks, results for high-connectivity networks may be confounded, leading to missing edges. Several personalized federated learning methods have been proposed to address such issues. For example, FedNova [58] normalizes aggregated updates to mitigate the impact of heterogeneous local computational workloads. More importantly, asynchronous techniques are particularly important for heterogeneous cases. Even though our asynchronous technique does not lead to a loss of accuracy, applying asynchronous methods remains crucial and represents a promising research direction, especially in the context of reinforcement learning, as explored by Espeholt et al. [14].